Project

General

Profile

03192014 Minutes

Attending: Gavin, Mat, Kuldeep, Tian, Micheal, Dom

Jose - Reported watchdog whiteness.

We will add a bulletin to properly indicate how much white space is allow in the histogram during catch-up.

Andrew reported:
- It has been set to 7 days (i.e. files must be older than 7 days as well as meet all other archive criteria before they can be erased from disk at FD)

Nearline-
NUmI Trigger rate drop to 0.1 Hz at 11 am
Cosmics trigger rate also low.

Hits/spill is caused by dew point - possibly influence tracks.

Kuldeep -
Working on understanding how slicer is responding to noisy detector conditions.
She has files. Will up date on Friday.

Tian -
Jon Paley - interest in min adc versus time.
Min adc versus
Showing 10936.

Michael -
Onmon ND files are being produced at some level.

Gavin -
Is freaking out... ND is going to APD inside of 2 weeks.